OpenStudy (anonymous):

Solve the problem. Make sure your answer is in simplest form and improper fractions are expressed as mixed numbers. 3(8/9) x 4(4/5)=?

OpenStudy (bamboo888):

I'll help

OpenStudy (karlaltr):

All you must do is make the equation as small as possible by using bedmas!:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay there is not much explaining to do. All you have to do is solve 3(8/9) get the answer and then multiply it with the answer from 4(4/5). When you do this you should get: \[\frac{ 128 }{ 15 }\] OR \[8 \frac{ 8 }{ 15 }\]

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @narlystarly

OpenStudy (bamboo888):

18 2/3

OpenStudy (anonymous):

thank you so much

OpenStudy (bamboo888):

\[18\frac{ 2 }{ 3 }\] Because when you multiply or divide a mixed number you must turn it into a improper fraction first.

OpenStudy (anonymous):

its needs to be a mixed number

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (bamboo888):

How you do it is by: ***multiplying the denominator by the whole number then adding the numerator.***

OpenStudy (bamboo888):

Yes.

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

Solve the problem. Make sure your answer is in simplest form and improper fractions are expressed as mixed numbers. A packer fills 7(1/3) boxes in 1 hour. How many boxes can the packer fill in 8(1/4) hours?

OpenStudy (bamboo888):

Early cross simplifying is a lot easier @Tom_Boy_Rebel

OpenStudy (bamboo888):

8/9

OpenStudy (anonymous):

it needs to be a mixed number

OpenStudy (bamboo888):

???

OpenStudy (bamboo888):

so it's 7 1/3 divided by 8 1/4 right?

OpenStudy (anonymous):

yes

OpenStudy (bamboo888):

i tried 5 times and still got 8/9

OpenStudy (anonymous):

ok i will see what i can do but thanks for trying and helping me and here

OpenStudy (bamboo888):

Thank you and you're welcome. ;)